# Qingdao 266237, China

### v1

@@ 2.2 Efficient Batch ssPMT
To improve the efficiency of mq-ssPMT, we abstract a new functionality called batch ssPMT, which is essentially the batched version of single-query ssPMT. In the batch ssPMT functionality, the sender $\mathcal{S}$ inputs $n$ disjoint sets $X_{1},\cdots,X_{n}$, and the receiver $\mathcal{R}$ inputs $n$ elements $y_{1},\cdots,y_{n}$. $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, whose $i$th bit is 1 if $y_{i}\in X_{i}$, otherwise 0. The relationship between batch ssPMT and mq-ssPMT is two-fold: In terms of functionality, batch ssPMT is to mq-ssPMT what batch OPRF is to multi-point OPRF, testing a batch of elements’ memberships across distinct sets rather than a common set; In terms of efficiency, batch ssPMT admits a more efficient construction, providing a superior alternative to mq-ssPMT in the context of MPSU.
We adopt the following two-step approach to build the batch ssPMT protocol: (1) $\mathcal{S}$ and $\mathcal{R}$ invoke batched oblivious programmable pseudorandom function (batch OPPRF) [KMP+17, PSTY19]. To elaborate, for the $i$th ssPMT instance ($i=1,\cdots,n$), $\mathcal{S}$ choose a random $s_{i}$, and creates a set $E_{i}$ comprising key-value pairs with elements from $X_{i}$ as keys and $s_{i}$ as all elements’ corresponding values. Then $\mathcal{S}$ and $\mathcal{R}$ invoke batch OPPRF of size $n$, where in the $i$th instance of OPPRF, $\mathcal{S}$ inputs $E_{i}$, while $\mathcal{R}$ inputs $y_{i}$ and receives $t_{i}$. According to the definition of OPPRF (cf. Section 3.5), if $y_{i}\in X_{i}$, $t_{i}$ equals $s_{i}$, otherwise $t_{i}$ is pseudorandom to $s_{i}$. If $y_{i}\notin X_{i}$, the probability of $t_{i}=s_{i}$ is $2^{-\gamma}$, where $\gamma$ is the output length of OPPRF, i.e., the bit length of $s_{i}$ and $t_{i}$. To ensure that the probability of any $t_{i}\neq s_{i}$ occurring is less than $2^{-\sigma}$ ($\sigma$ is the statistical security parameter), we set $\gamma\geq\sigma+\log n$ so that $n\cdot 2^{-\gamma}\leq 2^{-\sigma}$. (2) $\mathcal{S}$ and $\mathcal{R}$ invoke $n$ instances of secret-shared private equality test (ssPEQT) [PSTY19, CGS22], where in the $i$th instance of ssPEQT, $\mathcal{S}$ inputs $s_{i}$ and $\mathcal{R}$ inputs $t_{i}$. If $s_{i}=t_{i}$, they receive secret shares of 1, otherwise 0. It’s easy to verify that if $y_{i}\in X_{i}$, $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1, otherwise they receive shares of 0 with overwhelming probability.
@@
Now we start to build an alternative to mq-ssPMT using our batch ssPMT. Recall that in mq-ssPMT, the sender $\mathcal{S}$’s input is a single set $X$. First, $\mathcal{S}$ and $\mathcal{R}$ preprocess their inputs through hashing to bin technique: $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign its input elements $y_{1},\cdots,y_{n}$ to $B$ bins via Cuckoo hashing [PR04], which ensures that each bin accommodates at most one element. At the same time, $\mathcal{S}$ assigns each $x\in X$ to all bins determined by $h_{1}(x),h_{2}(x),h_{3}(x)$. Then the parties invoke $B$ instances of ssPMT, where in the $j$th instance, $\mathcal{S}$ inputs the subset $X_{j}\in X$ containing all elements mapped into its bin $j$ and $\mathcal{R}$ inputs the sole element mapped into its bin $j$. If some $y_{i}$ is mapped to bin $j$, and $y_{i}\in X$, then $\mathcal{S}$ certainly maps $y_{i}$ to bin $j$ (and other bins) as well, i.e., $y_{i}\in X_{j}$. Therefore, for each bin $j$ of $\mathcal{R}$, if the inside element $y_{i}$ is in $X$, we have $y_{i}\in X_{j}$, then $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1 from the $i$th instance of batch ssPMT. Otherwise, we have $y_{i}\notin X_{j}$, then $\mathcal{S}$ and $\mathcal{R}$ receive shares of 0.
@@
The above process in fact realizes a different functionality with mq-ssPMT. In the above construction, the query sequence of $\mathcal{R}$ is determined by the Cuckoo hash positions of its input elements. Since the Cuckoo hash positions depends on the whole input set of $\mathcal{R}$, and all shares is arranged in the order of the parties’ Cuckoo hash positions, the reconstruction stage may leak some information about the parties’ input sets to $P_{1}$ and hence cannot be simulated. Therefore, it is crucial to eliminate the dependence of shares’ order on the parties’ Cuckoo hash positions before the reconstruction, while fortunately, it has been realized by the execution of multi-party secret-shared shuffle as we discussed before. For the difference in efficiency, refer to Section 7.1.
@@ 2.3 SK-MPSU from Batch ssPMT and mss-ROT
By replacing mq-ssPMT with the construction using batch ssPMT plus hash to bin technique as previous, we can improve the performance of LG protocol significantly. However, the improved protocol still relies on the non-collusion assumption. So it remains to show that how to eliminate the non-collusion assumption to obtain a SK-MPSU protocol in the standard semi-honest model.
@@
First, let us figure out that why LG protocol needs the non-collusion assumption. We still use the aforementioned three-party PSU protocol as illustration: After the secret sharing process of $P_{3}$, all parties hold the secret shares of
$\begin{cases}x\|\mathsf{H}(x)&x\in Y_{3}\\ r_{3,1}^{0}\oplus r_{3,1}^{1}\oplus x\|\mathsf{H}(x)&x\in Y_{1}\\ r_{3,2}^{0}\oplus r_{3,2}^{1}\oplus x\|\mathsf{H}(x)&x\in Y_{2}\\ r_{3,1}^{1}\oplus r_{3,2}^{1}\oplus r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)&x\in X_{2}\cap X_{1}\end{cases}$
@@
In the final stage, if $x\in Y_{3}$, $P_{1}$ reconstructs the secret $x\|\mathsf{H}(x)$. Otherwise, the secret is uniformly at random in the $P_{1}$’s view so that $P_{1}$ gains no information about $x$. Nevertheless, this is only true when $P_{1}$ does not collude with $P_{3}$. Since there is no randomness in $P_{3}$’s view (In the two invocations of ROT, $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$ and $r_{3,2}^{0},r_{3,2}^{1}$), for each $x\in X_{3}$, the coalition of $P_{1}$ and $P_{3}$ can easily check the reconstructed values, and distinguish the cases that $x\in Y_{1}$, $x\in Y_{2}$, and $x\in X_{2}\cap X_{1}$, which reveals the information of $P_{2}$’s inputs.
@@
To resolve the above security problem, we generalize ROT into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT). Suppose there are $d<m$ involved parties and two of them, denoted $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$, possessing choice bits $b_{0}$ and $b_{1}$ respectively. The mss-ROT functionality gives each involved party $P_{i}$ ($i\in\{1,\cdots,d\}$) two outputs $r_{i}$ and $\Delta_{i}$, such that if $b_{0}\oplus b_{1}=0$, $r_{1}\oplus\cdots\oplus r_{d}=0$. Otherwise, $r_{1}\oplus\cdots\oplus r_{d}=\Delta_{1}\oplus\cdots\oplus\Delta_{d}$. Note that if $b_{0}\oplus b_{1}=1$, the parties share a value which is uniformly random in the view of a coalition of any $d-1$ parties.
@@
Equipped with this new primitive, the above protocol can be amended to achieve security under arbitrary collusion: After $P_{3}$ and $P_{1}$ receiving shares $e_{3,1}$ and $e_{1,3}$, while $P_{3}$ and $P_{2}$ receiving shares $e_{3,2}$ and $e_{2,3}$, $P_{1},P_{2},P_{3}$ invoke mss-ROT twice. In the first invocation, $P_{3}$ and $P_{1}$ act as $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ holding as choice bits $e_{3,1}$ and $e_{1,3}$. $P_{1},P_{2},P_{3}$ receives $r_{1,31},r_{2,31},r_{3,31}$ and $\Delta_{1,31},\Delta_{2,31},\Delta_{3,31}$ separately. If $e_{3,1}\oplus e_{1,3}=0$, $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=0$, otherwise $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=\Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}$. Likewise, in the second invocation, $P_{3}$ and $P_{2}$ hold as choice bits $e_{3,2}$ and $e_{2,3}$. $P_{1},P_{2},P_{3}$ receives $r_{1,32},r_{2,32},r_{3,32}$ and $\Delta_{1,32},\Delta_{2,32},\Delta_{3,32}$ separately. If $e_{3,2}\oplus e_{2,3}=0$, $r_{1,32}\oplus r_{2,32}\oplus r_{3,32}=0$, otherwise $r_{1,32}\oplus r_{2,32}\oplus r_{3,32}=\Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}$. Finally, $P_{1}$ sets its share $sh_{1}$ to be $r_{1,31}^{0}\oplus r_{1,32}^{0}$. $P_{2}$ sets its share $sh_{2}$ to be $r_{2,31}^{0}\oplus r_{2,32}^{0}$. $P_{3}$ sets its share $sh_{3}$ to be $r_{3,31}^{0}\oplus r_{3,32}^{0}\oplus x\|\mathsf{H}(x)$.
@@
$\begin{cases}x\|\mathsf{H}(x)&x\in Y_{3}\\ \Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}\oplus x\|\mathsf{H}(x)&x\in Y_{1}\\ \Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in Y_{2}\\ \Delta_{1,31}\oplus\Delta_{2,31}\oplus\Delta_{3,31}\oplus\Delta_{1,32}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in X_{2}\cap X_{1}\end{cases}$
It is immediate that if $x\notin Y_{3}$, the secret-shared value is uniform and independent of $x$ in the view of a coalition of any two parties. Hence, the reconstruction reveals no additional information to $P_{1}$ even if it colludes with $P_{2}$ or $P_{3}$.
@@ 2.4 PK-MPSU from Batch ssPMT and MKR-PKE
As previously mentioned, there are no existing MPSU works that achieve both linear computation and communication complexity. Thereinto, LG protocol and our SK-MPSU fail to achieve linear complexity because in the secret-sharing based framework, each party holds shares of $(m-1)n$ elements and in the final reconstruction stage, all parties except $P_{1}$ send their shares to $P_{1}$, then $P_{1}$ reconstruct $(m-1)n$ secrets to pick up all elements comprising the union. It is clear that even in this single stage, the computation and communication complexity of $P_{1}$ deviate from the linear criteria. So we have to turn to another approach.
We notice that the most advanced work in terms of asymptotic complexity is [GNT23], achieving near-linear complexity. We start by analyzing [GNT23]: The first phase enables $P_{1}$ to somehow acquire encrypted $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$ for $2\leq i\leq m$, interspersed with encrypted dummy messages filling positions of the duplicate elements. Note that if $P_{1}$ decrypts these ciphertexts by itself, then it can establish a mapping associating each element $x\in X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$ with the party to whom $x$ belongs. To address this problem, they introduce a PKE variant, MKR-PKE, with several tailored properties, enabling all parties to conduct a collaborative decryption and shuffle procedure.
@@
The second phase, which is the aforemetioned collaborative decryption and shuffle procedure, works as follows: $P_{1}$ sends the ciphertexts to $P_{2}$ after permuting them using a random permutation $\pi_{1}$. $P_{2}$ performs partial decryption on the received ciphertexts before rerandomization, and permutes them using a random permutation $\pi_{2}$. Then it forwards the permuted partially decrypted ciphertexts to $P_{3}$. This iterative process continues until the last party, $P_{m}$, who returns its permuted partially decrypted ciphertexts to $P_{1}$. $P_{1}$ decrypts the ultimate ciphertexts, filters out the dummy elements, retains the desired set $X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$, and appends the elements in $X_{1}$, to obtain the union.
We identify that their non-linear complexity stems from their construction of the first phase. Despite that they also use hashing to bin technique, their construction does not support amortizing the batching cost. In the case $B=O(n)$ bins are considered, each party has to execute mOT and cOPRF of size $O(\log\log n)$ per bin, resulting in overall $O(n\log\log n)$ computation and communication complexity. Besides, their insecurity against arbitrary collusion roots in the construction of the first phase (cf. Appendix 0.A). In contrast, the second phase already has linear computation and communication complexity, as well as security against arbitrary collusion. [^5] As we have said, the second phase addresses the same problem solved by the multi-party secret-shared shuffle, meanwhile there currently exists no construction for multi-party secret-shared shuffle that achieves total linear complexity. This is the root why our PK-MPSU protocol can achieve total linear complexity while our SK-MPSU protocol cannot. The key distinction lies in our PK-MPSU protocol’s departure from the secret-shared paradigm, opting instead for a PKE-based approach to replace the non-linear component. Therefore, the task of devising linear-complexity MPSU reduces to devising a linear-complexity and secure construction against arbitrary collusion for the first phase.
Our first attempt is as follows: For $2\leq i\leq m$, each $P_{i}$ and $P_{1}$ execute batch ssPMT (after preprocessing their inputs using hashing to bin, we ignore this for the sake of simplicity in the later explanation) and receive secret shares as the choice bits for the subsequent execution of two-choice-bit OT [^6] The difference of functionality between two-choice-bit OT and two-choice-bit ROT is that $\mathcal{S}$ inputs two messages $m_{0},m_{1}$ instead of obtaining two uniform messages., where for each $x\in X_{i}$, $P_{i}$ acts as $\mathcal{S}$ and inputs the encrypted $x$ and an encrypted dummy message. As a result, if $x$ belongs to the difference set $X_{i}\setminus X_{1}$, $P_{1}$ receives the encrypted $x$; otherwise, it receives the encrypted dummy message. Nevertheless, the goal of the first phase is to obtain the encrypted difference set $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$, which is the subset of $X_{i}\setminus X_{1}$.
In order to attain the above goal, our next strategy is to let each $P_{i}$ engage in a procedure with each $P_{j}$ ($1<j<i$) before executing the batch ssPMT and two-choice-bit OT with $P_{1}$, so that $P_{j}$ can “obliviously” replace the encrypted items in the intersection $X_{j}\cap X_{i}$ with encrypted dummy messages. The procedure unfolds as follows.
@@
$P_{i}$ plays the role of $\mathcal{R}$ and executes the batch ssPMT with $P_{j}$. For each $x\in X_{i}$, $P_{i}$ and $P_{j}$ receive shares $e_{i,j}$ and $e_{j,i}$. $P_{i}$ initializes a message $m_{0}$ to the encrypted $x$ and $m_{1}$ to an encrypted dummy message, then executes two-choice-bit OT with $P_{j}$. $P_{i}$ and $P_{j}$ input $e_{i,j}$ and $e_{j,i}$ as their choice bits. $P_{i}$ acts as $\mathcal{S}$ and inputs $m_{0}$ and $m_{1}$. $P_{j}$ acts as $\mathcal{R}$ and receives $\mathsf{ct}=m_{e_{i,j}\oplus e_{j,i}}$. If $x\notin X_{j}$, $\mathsf{ct}$ is the ciphertext of $x$; otherwise, the ciphertext of dummy message. $P_{j}$ rerandomizes $\mathsf{ct}$ to $\mathsf{ct}^{\prime}$ and returns $\mathsf{ct}^{\prime}$ to $P_{i}$. $P_{i}$ rerandomizes $\mathsf{ct}^{\prime}$, then updates $m_{0}$ to $\mathsf{ct}^{\prime}$ and $m_{1}$ to the rerandomized $m_{1}$ before repeating the above procedure with the next party $P_{j+1}$. The final values of $m_{0}$ and $m_{1}$ shall be used during the subsequent invocation of two-choice-bit OT with $P_{1}$.

### v4

@@ 2.2 Efficient Batch ssPMT
    To improve the efficiency of mq-ssPMT, we introduce a new functionality called batch ssPMT: The sender $\mathcal{S}$ inputs $n$ disjoint sets $X_{1},\cdots,X_{n}$, and the receiver $\mathcal{R}$ inputs $n$ elements $y_{1},\cdots,y_{n}$. $\mathcal{S}$ and $\mathcal{R}$ receive secret shares of a bit vector of size $n$, where the $i$th bit is 1 if $y_{i}\in X_{i}$, and 0 otherwise.
    Compared to mq-ssPMT, batch ssPMT enables the testing of element membership across multiple distinct sets rather than within a single common set. Their relationship is analogous to that between batched oblivious pseudorandom function (batch OPRF) [KKRT16] and multi-point oblivious pseudorandom function (multi-point OPRF) [PRTY19, CM20]. Thanks to its batching nature, batch ssPMT admits a much more efficient construction, lending it a superior alternative to mq-ssPMT in the context of MPSU, by coupling with hashing to bins.
@@
    The alternative to mq-ssPMT works as follows: First, $\mathcal{S}$ and $\mathcal{R}$ preprocess inputs through hashing to bins technique. $\mathcal{R}$ uses hash functions $h_{1},h_{2},h_{3}$ to assign $y_{1},\cdots,y_{n}$ to $B$ bins via Cuckoo hashing [PR04], ensuring that each bin accommodates at most one element. At the same time, $\mathcal{S}$ assigns each $x\in X$ to all bins determined by $h_{1}(x),h_{2}(x),h_{3}(x)$. Then they invoke $B$ instances of ssPMT, where in the $j$th instance, $\mathcal{S}$ inputs the subset $X_{j}\in X$ containing all elements mapped into its bin $j$, while $\mathcal{R}$ inputs the sole element mapped into its bin $j$. If some $y_{i}$ is mapped to bin $j$ and $y_{i}\in X$, then $\mathcal{S}$ certainly maps $y_{i}$ to bin $j$ (and other bins) as well, ensuring $y_{i}\in X_{j}$. Therefore, for each bin $j$ of $\mathcal{R}$, if the inside element $y_{i}$ is in $X$, $y_{i}\in X_{j}$, $\mathcal{S}$ and $\mathcal{R}$ receive shares of 1 from the $i$th instance of batch ssPMT. Otherwise, $y_{i}\notin X_{j}$, $\mathcal{S}$ and $\mathcal{R}$ receive shares of 0.
@@
    The functionality realized by the above construction differs slightly from mq-ssPMT, as the query sequence of $\mathcal{R}$ is determined by the Cuckoo hash positions of its input elements. Since the Cuckoo hash positions depends on the entire input set of $\mathcal{R}$, and all shares is arranged according to the parties’ Cuckoo hash positions, a straightforward reconstruction may leak information about the parties’ input sets to $P_{1}$. Therefore, we have to eliminate the dependence of shares’ order on the parties’ Cuckoo hash positions during the reconstruction phase. Our crucial insight is that the multi-party secret-shared shuffle used in the reconstruction phase eliminates not only the correspondence between secret shares and difference sets, but also this dependence on the hash positions. As a result, our construction can plug in the secret-sharing based framework seamlessly, without introducing any information leakage or additional overhead.
    Unlike the mq-ssPMT, which heavily relies on general 2PC, our batch ssPMT is built on two highly efficient primitives, batched oblivious programmable pseudorandom function (batch OPPRF) [KMP+17, PSTY19] and secret-shared private equality test (ssPEQT) [PSTY19, CGS22], where the former has an extremely fast specialized instantiation through vector oblivious linear evaluation (VOLE), while the later is composed of a small circuit, allowing for efficient implementation with general 2PC. Therefore, our batch ssPMT greatly reduces dependency on general 2PC, leading to a substantial performance improvement. For more details, refer to Appendix 0.E.2.2.
@@ 2.3 SK-MPSU from Batch ssPMT and mss-ROT
    By replacing mq-ssPMT with the construction in the previous section, we improve LG’s performance significantly. However, the resulting protocol still relies on the non-collusion assumption. We proceed to show how to eliminate it.
@@
    First, we explain why LG requires the non-collusion assumption using the three-party example: After the secret-sharing process of $P_{3}$, all parties hold the secret shares of
    $\begin{cases}x\|\mathsf{H}(x)&x\in X_{3}\setminus(X_{1}\cup X_{2})\\ r_{3,1}^{0}\oplus r_{3,1}^{1}\oplus x\|\mathsf{H}(x)&x\in(X_{2}\cup X_{3})\setminus X_{1}\\ r_{3,2}^{0}\oplus r_{3,2}^{1}\oplus x\|\mathsf{H}(x)&x\in(X_{1}\cup X_{3})\setminus X_{2}\\ r_{3,1}^{1}\oplus r_{3,2}^{1}\oplus r_{3,1}^{0}\oplus r_{3,2}^{0}\oplus x\|\mathsf{H}(x)&x\in X_{1}\cap X_{2}\cap X_{3}\end{cases}$
@@
    If $x\in Y_{3}$, $P_{1}$ reconstructs the secret $x\|\mathsf{H}(x)$. If $x\notin Y_{3}$, the secret is random so that $P_{1}$ gains no information about $x$. Nevertheless, this is only guaranteed when $P_{1}$ does not collude with $P_{3}$. Recall that $P_{3}$ receives $r_{3,1}^{0},r_{3,1}^{1}$ and $r_{3,2}^{0},r_{3,2}^{1}$ from the ROT execution with $P_{1}$ and $P_{2}$ respectively, so the coalition of $P_{1}$ and $P_{3}$ can easily check the secret, and distinguish the distinct cases, which reveals information of $P_{2}$’s inputs.
@@
    To fix this security issue, we generalize ROT into multi-party setting, which we call multi-party secret-shared random oblivious transfer (mss-ROT). Consider $d$ involved parties, with two parties, denoted as $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$, possessing choice bits $b_{0}$ and $b_{1}$ respectively. There is a subset $J$ of $\{1,\cdots,d\}$ so that each party $P_{j}$ ($j\in J$) holds $\Delta_{j}$ as input. The mss-ROT functionality outputs a random $r_{i}$ to each party $P_{i}$, s.t. if $b_{0}\oplus b_{1}=0$, $\bigoplus_{i=1}^{d}r_{i}=0$. Otherwise, $\bigoplus_{i=1}^{d}r_{i}=\bigoplus_{j\in J}\Delta_{j}$.
@@
    Equipped with this new primitive, the secret-sharing process of $P_{3}$ is revised as follows: After $P_{3}$ and $P_{1}$ receiving shares $e_{3,1}$ and $e_{1,3}$, $\{P_{1},P_{2},P_{3}\}$ invoke mss-ROT, where $P_{3}$ and $P_{1}$ act as $P_{\text{ch}_{0}}$ and $P_{\text{ch}_{1}}$ holding $e_{3,1}$ and $e_{1,3}$, while $P_{2}$ and $P_{3}$ samples $\Delta_{2,31},\Delta_{3,31}$ uniformly as inputs. $P_{1},P_{2},P_{3}$ receives $r_{1,31},r_{2,31},r_{3,31}$ separately. If $e_{3,1}\oplus e_{1,3}=0$, $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=0$, otherwise $r_{1,31}\oplus r_{2,31}\oplus r_{3,31}=\Delta_{2,31}\oplus\Delta_{3,31}$. After $P_{3}$ and $P_{2}$ receiving shares $e_{3,2}$ and $e_{2,3}$, $\{P_{2},P_{3}\}$ invoke mss-ROT, where $P_{3}$ and $P_{2}$ hold choice bits $e_{3,2}$ and $e_{2,3}$, while $P_{2}$ and $P_{3}$ samples $\Delta_{2,32},\Delta_{3,32}$ uniformly as inputs. $P_{2}$ and $P_{3}$ receives $r_{2,32}$ and $r_{3,32}$ separately. If $e_{3,2}\oplus e_{2,3}=0$, $r_{2,32}\oplus r_{3,32}=0$, otherwise $r_{2,32}\oplus r_{3,32}=\Delta_{2,32}\oplus\Delta_{3,32}$. Finally, $P_{1}$ sets $s_{3,1}=r_{1,31}$. $P_{2}$ sets $s_{3,2}=r_{2,31}\oplus r_{2,32}$. $P_{3}$ sets $s_{3,3}=r_{3,31}\oplus r_{3,32}\oplus x\|\mathsf{H}(x)$.
@@
    $\begin{cases}x\|\mathsf{H}(x)&x\in X_{3}\setminus(X_{1}\cup X_{2})\\ \Delta_{2,31}\oplus\Delta_{3,31}\oplus x\|\mathsf{H}(x)&x\in(X_{2}\cup X_{3})\setminus X_{1}\\ \Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in(X_{1}\cup X_{3})\setminus X_{2}\\ \Delta_{2,31}\oplus\Delta_{3,31}\oplus\Delta_{2,32}\oplus\Delta_{3,32}\oplus x\|\mathsf{H}(x)&x\in X_{1}\cap X_{2}\cap X_{3}\end{cases}$
    It is immediate that if $x\notin Y_{3}$, the secret is random in the view of the coalition of $\{P_{1},P_{2}\}$ or $\{P_{1},P_{3}\}$. Since the coalition of $\{P_{2},P_{3}\}$ cannot know the secret, the protocol achieves security under arbitrary collusion.
@@ 2.4 PK-MPSU from Batch ssPMT and MKR-PKE
    There are no existing MPSU protocols with linear computation and communication complexity. Notably, it is impossible to achieve this goal within the secret-sharing based framework, given that $P_{1}$ must reconstruct $(m-1)n$ secrets with each secret comprising $m$ shares. Therefore, we have to seek for another technical route.
@@
    Our start point is the existing work [GNT23] with the best asymptotic complexity $mn(\log n/\log\log n)$. We distill their protocol into a framework with two phases: The first phase allows $P_{1}$ to somehow obtain encrypted $X_{i}\setminus(X_{1}\cup\cdots\cup X_{i-1})$ for $2\leq i\leq m$, interspersed with encrypted dummy messages filling the positions of duplicate elements. However, if $P_{1}$ decrypts ciphertexts by itself, then it could associate each element in $X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$ with the specific party it originates from. This problem is addressed by the second phase, which is the collaborative decryption and shuffle procedure, making use of a PKE variant MKR-PKE: $P_{1}$ sends the ciphertexts to $P_{2}$ after permuting them using a random permutation $\pi_{1}$. $P_{2}$ performs a partial decryption on the received ciphertexts, rerandomizes them, permutes them with a random permutation $\pi_{2}$, and forwards the ciphertexts to $P_{3}$. This iterative process continues until the last party, $P_{m}$, returns its permuted partially decrypted ciphertexts to $P_{1}$. Finally, $P_{1}$ fully decrypts the ciphertexts, filters out the dummy elements, retains the set $X_{1}\cup\cdots\cup X_{m}\setminus X_{1}$, and appends $X_{1}$ to compute the union.
    We identify that both the non-linear complexities and insecurity against arbitrary collusion of [GNT23] arise from their construction in the first phase. Therefore, the task of building a linear-complexity MPSU reduces to devising a linear-complexity and secure construction for this phase.
    A rough idea is as follows: Let $P_{i}$ hold a set of encrypted elements in $X_{i}$. For $2\leq i\leq m$, each $P_{i}$ engage in a procedure with each $P_{j}$ ($1<j<i$) such that $P_{j}$ can “obliviously” replace the encrypted elements in $X_{i}\cap X_{j}$ with encrypted dummy messages. After the procedure, $P_{i}$ holds a set of encrypted elements in $X_{i}\setminus(X_{2}\cup\cdots\cup X_{i-1})$. The replacement of encrypted elements in $X_{1}$ is handled at the end.
@@
    The replacing procedure between each $P_{i}$ and $P_{j}$: (1) $P_{i}$ acts as $\mathcal{R}$ and executes batch ssPMT (after preprocessing their inputs using hashing to bins, which we omit here for simplicity) with $P_{j}$. For each $x\in X_{i}$, $P_{i}$ and $P_{j}$ receive shares $e_{i,j}$ and $e_{j,i}$. (2) If $j=2$, $P_{i}$ initializes $m_{0}$ as the encrypted $x$ and $m_{1}$ as an encrypted dummy message. (3) $P_{i}$ acts as $\mathcal{S}$ and executes two-choice-bit OT [^7] The difference from two-choice-bit ROT is that $\mathcal{S}$ inputs two specific messages $m_{0},m_{1}$ instead of obtaining two uniform messages. with $P_{j}$, where $P_{i}$ and $P_{j}$ input $e_{i,j}$ and $e_{j,i}$ as choice bits. $P_{i}$ inputs $m_{0},m_{1}$ and $P_{j}$ receives $\mathsf{ct}=m_{e_{i,j}\oplus e_{j,i}}$. If $x\notin X_{j}$, $\mathsf{ct}$ is the ciphertext of $x$; otherwise, the ciphertext of dummy message. (4) $P_{j}$ rerandomizes $\mathsf{ct}$ to $\mathsf{ct}^{\prime}$ and returns $\mathsf{ct}^{\prime}$ to $P_{i}$. (5) $P_{i}$ rerandomizes $\mathsf{ct}^{\prime}$, updates $m_{0}$ to $\mathsf{ct}^{\prime}$, and rerandomizes $m_{1}$ before repeating the procedure with the next party $P_{j+1}$. Note that only if all conditions of $x\notin X_{j}$ are satisfied, the final $m_{0}$ is an encrypted message of $x$.
    For $2\leq i\leq m$, each $P_{i}$ and $P_{1}$ execute batch ssPMT and receive secret shares as choice bits for the subsequent execution of two-choice-bit OT. For each $x\in X_{i}$, $P_{i}$ acts as $\mathcal{S}$ and inputs the final $m_{0}$ of the above procedure and an encrypted dummy message. As a result, if $x\notin X_{1}$, $P_{1}$ receives the final $m_{0}$; otherwise, it receives the encrypted dummy message. This excludes all elements in $X_{1}$ from the encrypted set $X_{i}\setminus(X_{2}\cup\cdots\cup X_{i-1})$ and lets $P_{1}$ obtain the results.
    The linear complexities of our construction are attributed to two key facts: First, this framework ensures that each interaction between $P_{i}$ and $P_{j}$ only involves the input sets of themselves. Second, our batch ssPMT protocol has linear computation and communication complexity (cf. Section 4).